# Are We Asking the Right Questions?: Designing for Community Stakeholders’ Interactions with AI in Policing

### 08 Feb 2024 v1

@@ 5.2. AI Applications as Boundary Objects and Educational Tools
Through our observations of lab activities conducted with community members, we learned that algorithmic crime mapping served as an educational tool that taught them several key aspects of algorithmic decision-making. First, the application focused on familiar city neighborhoods which helped establish a degree of comfort in the lab activity such that community members were not overwhelmed and were able to freely deliberate. Second, by engaging with parameters, community members understood that the underlying algorithm was simply generating estimates (i.e., hotspots) and not some objective truth. Their choice of parameters (i.e., size, smoothness, and distance between hotspots) impacted how the results were generated on the map. Even though we had given the participants a brief overview of each parameter, they developed a more nuanced understanding by experimenting with them (e.g., the degree of numeric change causing the degree of visual change on the map). Participants also found the results to be more interpretable because they were able to ground these empirical findings within their local and contextual knowledge of different neighborhoods. Several community members and technical experts unexpectedly sought to compare different maps simultaneously, causing frustration for some participants. However, a few quickly addressed the issue by using the ‘snipping tool’ to capture and preserve the current map before generating a new one. As participants engaged with the maps and understood the underlying data, they started to raise concerns regarding issues such as speed traps, racial profiling in low-income areas, and policing disparities between suburban and inner-city regions. Here, the application acted as a boundary object that facilitated participants in brainstorming issues and initiating discussions with the researchers. However, some participants found the tool useful if operated in the public interest, envisioning it as a resource to learn about different neighborhoods. For example, a participant suggested using the tool to decide whether to purchase garage parking in areas prone to car thefts. By engaging and deliberating with the application, participants were able to develop an informed mistrust of the AI tool, i.e. - they were able to think of use cases where the AI can be used responsibly and inform decision-making but also articulated scenarios where the application could do harm (e.g., racial profiling). In sum, this study provides further evidence that integrating user control into intuitive AI applications that leverage local and contextual knowledge can effectively anchor AI concepts and educate community members. This is particularly significant given the expanding research focus on public interest technology (stapleton2022has, ; aragon2020civic, ; mccord2023beyond, ), community-driven AI (hadley2022covid, ; kuo2023understanding, ; stapleton2023seeing, ), and expanding participation in AI design (delgado2023participatory, ; zhang2010detecting, ; kuo2023understanding, ) within the SIGCHI community.
@@
On the other hand, our observations of domain experts (i.e., LEAs) revealed a more complicated story. LEAs demonstrated a tendency to anchor their street-level knowledge to specific landmarks and neighborhoods in the city, persistently sticking to the initial parameters despite the researchers urging them to explore different maps. This indicated anchoring bias, a cognitive bias where individuals rely too heavily on the first piece of information they encounter (the ”anchor”). Anchoring bias has been observed in domain experts across various disciplines (furnham2011literature, ; azzopardi2021cognitive, ; saposnik2016cognitive, ). However, as noted in Section 4.3.3, LEAs observed that the application lacked essential features for deriving meaningful information. As a result, they engaged with it at a higher level, using it to validate street-level knowledge and generate evidence for requesting specialized patrol zones. Our findings provide further evidence that AI systems that are meant to augment workers’ social decision-making (kawakami2022care, ) must seek to complement their decision-making abilities and overcome their limitations (kawakami2023training, ; guerdan2023ground, ). That is, worker-centered design principles are fundamental for laying the groundwork for human-AI decision-making and evaluating whether human+AI decisions lead to improved outcomes. We delve deeper into this perspective in the next section and offer implications for human-AI decision-making.
@@ 5.3. Implications for Human-AI Decision-Making
We learned that domain expertise played a significant role in determining LEAs’ interactions with the application. LEA used their domain expertise at all three levels of engagement: understanding the system, interacting with the system, and interpreting the results; highlighting that their decision-making was a function of their domain (i.e., theoretical) knowledge where the application (i.e., empirical knowledge) was mostly used to anchor and validate pre-existing information. That is, empirical information did not augment the practitioners’ knowledge by introducing factors they would not have considered otherwise. This was evident from the LEAs choosing not to generate new maps as well as the results of the NASA-TLX survey that showed that they were mentally engaged but in doing so relied too heavily on the initial information presented to them which in turn influenced their subsequent decisions. This further suggests that there is a need for LEAs to unlearn their current AI practices and be trained to engage in critical use (kawakami2023training, ) - engaging in appropriate reliance by situating AI predictions against their contextual and domain knowledge.
@@
We also learned that target-construct mismatch (kawakami2023training, ; guerdan2023ground, ) can exist even in applications as intuitive as algorithmic crime mapping. Predicting hotspots across the city is supposed to help LEAs distribute and allocate resources more efficiently to deter crime (i.e., the target variable), however, we learned that LEAs were more interested in finding critical hyperlocal regions such as key intersections to create specialized patrol zones (i.e., construct of interest) where they could direct resources and also ensure the availability of backup for patrol cars. Here, as highlighted by Kawakami et al. (kawakami2023training, ), effective AI use depends on the practitioners’ ability to account for such misalignments.
@@
Here, practitioners can find themselves translating information from these two different sources which can often increase uncertainty and lead to unreliable decisions (kawakami2022improving, ; mcconvey2023human, ; saxena2022train, ). Systems that are not designed to complement the workers’ expertise (i.e., grounded in worker-centered design (fox2020worker, ; collective2022human, )) can further complicate human-AI decision-making and make it difficult to assess its efficacy. This argument may seem intuitive but needs to be clearly formulated. Domain experts are being asked to think in counter-intuitive ways and adopt empirical models that introduce data-driven insights to decision-making. However, without proper heuristic guidelines regarding how to incorporate these insights within decision-making such that the human+AI decision is an improvement over human decisions, such empirical models can either become a source of frustration for decision-makers (see for e.g., (cheng2022child, ; saxena2023rethinking, ; kawakami2022improving, ; watkins_art, )) or otherwise, as depicted in this case study, can be a source of anchoring bias. This was also evident from the LEA’s response who claimed that they were only looking for a satisfactory outcome from the system and not looking to derive new spatiotemporal information that would augment their decisions.
@@
Therefore, to be able to evaluate the efficacy of human-AI decision-making, it is first necessary to ensure that the AI complements the workers’ expertise and augments their decision-making in specific ways. For instance, LEAs shared that it would be meaningful to draw out the boundaries between districts (i.e., insert their hyperlocal contextual knowledge onto the map), look at the relationship between different types of crimes to assess any spatiotemporal relationships, or incorporate information derived from on-the-ground investigations. That is, worker-centered design of AI applications can help us better approach the evaluation of human+AI decisions by ensuring that AI complements workers’ expertise and augments their decisions in measurable ways. This common theme regarding human-AI decision-making has been observed across contexts (e.g., child welfare (kawakami2022improving, ; saxena2023algorithmic, ), nursing (watkins_art, ), hospitality (spektor2023designing, )) where workers draw attention to the nuances of their labor that the AI did not account for and which inadvertently augmented uncertainty and left the workers performing added labor to repair the disruption in organizational processes caused by the AI system.
@@
On the other hand, contextualizing empirical data points with qualitative narratives from police investigations can lead to a more robust AI system where explanations (i.e., a summary of key points) are generated alongside hotspots to help LEAs assess if new patterns of criminal activity are emerging in neighborhoods. Ehsan et al.’s (ehsan2023charting, ) framework for operationalizing explainable AI offers important practical considerations for designing for explanations where the nature of practice is highly collaborative.
@@ 5.4. Design Guidelines for Algorithmic Crime Mapping that Serves the Public Interest
Our study of a prominent algorithmic tool utilized in the criminal justice system found some comprehensive design principles for creating AI systems for public sector settings. However, as researchers, we must contend with the fact that most of the algorithmic tools used by law enforcement agents in the United States are privately developed with little room for comments from the public. In addition, considering the data being used by such technologies and their disparate use in urban areas for the policing of racial and low-income minorities, we believe that such technologies can lead to significant representational, interpersonal, and social algorithmic harms (shelby2023sociotechnical, ). Therefore, it was essential to explore the community members’ perspectives and gather their feedback. Community members drew upon their lived experiences, the technical participants used their data science expertise, and the LEAs leveraged their domain expertise.
@@
  • Technical participants discussed the elastic nature of data science practice and recommended a more collaborative process for decision-making to ensure more accountability and transparency around the practices of crime analysts. They also drew attention to the harmful feedback loop created by AI crime mapping.
@@
  • Worker-centered design of AI systems, as specified in the point above, can further make the system more transparent to workers by supporting AI explanations (ehsan2021expanding, ) that are grounded in the nature of practice.
  • It is important to recognize that a target-construct mismatch (kawakami2023training, ) is likely to exist, assess its impact on predicted outcomes, and whether the predicted target still aligns with the practitioners’ decision-making goals.
  • In response to algorithmic unfairness embedded in sociotechnical systems, significant attention has been paid to the biases embedded in datasets (e.g., disproportionate impact on some social groups). By putting the practitioners’ contextual knowledge into action, it is necessary to assess any process-oriented harms (i.e., harm to the decision-making process itself) that may occur (saxena2023algorithmic, ).
@@ 7. Conclusion
In this study, we conducted a mixed-methods investigation in a midwestern, mid-sized city in the United States. The goal of this study was to assess how community stakeholders from different backgrounds interacted with and interpreted a crime mapping algorithm. To achieve this, we designed an in-lab activity where participants were asked to interact with the tool, and subsequently, identify hotspots on the map. Additionally, we sought to gain insights into perspectives through semi-structured interviews conducted immediately after the activity. We learned that community members found utility in algorithmic crime mapping but were concerned about the incentive structures within police departments themselves that often racially profile and target low-income communities. Technical participants, on the other hand, drew attention to the elastic and subjective nature of data science practice and the need for a more collaborative and interdisciplinary process for decision-making. We learned that the LEAs interaction and interpretation of this application was significantly informed by their domain knowledge and they often used it as a means of validating their pre-existing street-level knowledge instead of trying to derive new information from it. We further show that AI systems need to be designed using worker-centered design principles if they are to augment human decisions such that human+AI decisions are an improvement over human-only and AI-only decisions.
@@ Appendix A ESTIMATING THE NUMBER OF HOTSPOTS
Fig 5 shows the distributions of minimum and maximum hotspots grouped by background. For our statistical analysis, we compared user-selected values to actual values for each map type (the solid line represents actual values). We conducted this analysis to ascertain whether people from different backgrounds tend to over or under-estimate hotspots.
@@ Appendix C Assessing mental workload of tasks
The third parameter we considered is performance (PF) which gauges workers’ success and satisfaction. In Figure 8, community members have medium PF scores (10-29), while technical participants and LEAs show somewhat high scores (30-49) (NASA, ). This suggests greater confidence in results for technical participants and LEAs compared to community members. We explain in Sections 4.1.1 and 4.1.2 that LEAs and technical participants are more familiar with such algorithmic systems than community members. Despite technical participants having data science expertise (average PF score of 45.6), LEAs, due to prior experience of working with algorithmic crime maps, exhibit similar confidence levels in their outputs.
The fourth parameter we considered is frustration level (FT), which assesses whether the job is frustrating or mentally demeaning such that the participant is left feeling insecure, desperate, offended, or disturbed when doing the task. The average FT is medium (10-29) for community members and low (0-9) for technical participants and LEAs. As we highlight in the qualitative results, community members were more frustrated because they wanted to generate and compare different maps. Notably, LEAs were less frustrated than technical participants, likely because they didn’t generate new maps, relying on the first map and connecting it with their street-level domain knowledge. Through this analysis, we can fairly say that the mental workload regarding the task is not very high which implies it is feasible to use such AI applications as boundary objects that can facilitate deeper conversations about algorithm design and use with a variety of stakeholders.

### 19 Mar 2024 v2

@@ 5.2. AI Applications as Boundary Objects and Educational Tools
Through our observations of lab activities conducted with community members, we learned that algorithmic crime mapping served as an educational tool that taught them several key aspects of algorithmic decision-making. First, the application focused on familiar city neighborhoods which helped establish a degree of comfort in the lab activity such that community members were not overwhelmed and were able to freely deliberate. Second, by engaging with parameters, community members understood that the underlying algorithm was simply generating estimates (i.e., hotspots) and not some objective truth. Their choice of parameters (i.e., size, smoothness, and distance between hotspots) impacted how the results were generated on the map. Even though we had given the participants a brief overview of each parameter, they developed a more nuanced understanding by experimenting with them (e.g., the degree of numeric change causing the degree of visual change on the map). Participants also found the results to be more interpretable because they were able to ground these empirical findings within their local and contextual knowledge of different neighborhoods.
Several community members and technical experts unexpectedly sought to compare different maps simultaneously, causing frustration for some participants. However, a few quickly addressed the issue by using the ‘snipping tool’ to capture and preserve the current map before generating a new one. As participants engaged with the maps and understood the underlying data, they started to raise concerns regarding issues such as speed traps, racial profiling in low-income areas, and policing disparities between suburban and inner-city regions. Here, the application acted as a boundary object that facilitated participants in brainstorming issues and initiating discussions with the researchers. However, some participants found the tool useful if operated in the public interest, envisioning it as a resource to learn about different neighborhoods. For example, a participant suggested using the tool to figure out if a neighborhood was prone to car thefts and subsequently decide whether to purchase garage parking. By engaging and deliberating with the application, participants were able to develop an informed mistrust of the AI tool, i.e. - they were able to think of use cases where the AI can be used responsibly and inform decision-making but also articulated scenarios where the application could do harm (e.g., racial profiling). In sum, this study provides further evidence that integrating user control into intuitive ADS that leverages local and contextual knowledge can effectively anchor AI concepts and educate community members. This is particularly significant given the expanding research focus on public interest technology (Stapleton et al., 2022b; Aragón et al., 2020; McCord and Becker, 2023), community-driven AI (Hadley et al., 2022; Kuo et al., 2023; Stapleton et al., 2023), and expanding participation in AI design (Delgado et al., 2023; Zhang et al., 2010; Kuo et al., 2023) within the SIGCHI community.
@@
On the other hand, our observations of domain experts (i.e., LEAs) revealed a more complicated story. LEAs demonstrated a tendency to anchor their street-level knowledge to specific landmarks and neighborhoods in the city, persistently sticking to the initial parameters despite the researchers urging them to explore different maps. This indicated anchoring bias, a cognitive bias where individuals rely too heavily on the first piece of information they encounter (the ”anchor”). Anchoring bias has been observed in domain experts across various disciplines (Furnham and Boo, 2011; Azzopardi, 2021; Saposnik et al., 2016). However, as noted in Section 4.3.3, LEAs shared that the application lacked essential features for deriving meaningful information. As a result, they engaged with it at a higher level, using it to validate street-level knowledge and generate evidence for requesting specialized patrol zones. Our findings provide further evidence that AI systems that are meant to augment workers’ social decision-making (Kawakami et al., 2022b) must seek to complement their decision-making abilities and overcome their limitations (Kawakami et al., 2023b; Guerdan et al., 2023). That is, worker-centered design principles are fundamental for laying the groundwork for human-AI decision-making and evaluating whether human+AI decisions lead to improved outcomes. We delve deeper into this perspective in the next section.
@@ 5.3. Implications for Human-AI Decision-Making
We learned that domain expertise played a significant role in determining LEAs’ interactions with the application. LEA used their domain expertise at all three levels of engagement: understanding the system, interacting with the system, and interpreting the results; highlighting that their decision-making was a function of their domain (i.e., theoretical) knowledge where the application (i.e., empirical knowledge) was mostly used to anchor and validate pre-existing information. That is, empirical information did not augment the practitioners’ knowledge by introducing factors they would not have considered otherwise. This was evident from the LEAs choosing not to generate new maps as well as the results of the NASA-TLX survey that showed that they were mentally engaged but in doing so relied too heavily on the initial information presented to them which in turn influenced their subsequent decisions. This further suggests that there is a need for LEAs to unlearn their current ADS practices and be trained to engage in critical use (Kawakami et al., 2023b) - engaging in appropriate reliance by situating AI predictions against their contextual and domain knowledge.
@@
We also learned that target-construct mismatch (Kawakami et al., 2023b; Guerdan et al., 2023) can exist even in applications as intuitive as algorithmic crime mapping. Predicting hotspots across the city is supposed to help LEAs distribute and allocate resources more efficiently to deter crime (i.e., the target variable), however, we learned that LEAs were more interested in finding critical hyperlocal regions such as key intersections to create specialized patrol zones (i.e., construct of interest) where they could direct resources and also ensure the availability of backup for patrol cars. Here, as highlighted by Kawakami et al. (Kawakami et al., 2023b), effective AI use depends on the practitioners’ ability to account for such misalignments.
@@
Here, practitioners can find themselves translating information from these two different sources which can often increase uncertainty and lead to unreliable decisions (Kawakami et al., 2022a; McConvey et al., 2023; Saxena et al., 2022b). Systems that are not designed to complement the workers’ expertise (i.e., grounded in worker-centered design (Fox et al., 2020; Collective et al., 2022)) can further complicate human-AI decision-making and make it difficult to assess its efficacy. This argument may seem intuitive but needs to be clearly formulated. Domain experts are being asked to think in counter-intuitive ways and adopt empirical models that introduce data-driven insights to decision-making. However, without proper heuristic guidelines regarding how to incorporate these insights within decision-making such that human+AI decisions are an improvement over human decisions, such empirical models can either become a source of frustration for decision-makers (see for e.g., (Cheng et al., 2022; Saxena et al., 2023; Kawakami et al., 2022a; Clare Elish and Anne Watkins, 2020)) or otherwise, as depicted in this case study, can be a source of anchoring bias. This was also evident from the LEA’s response who claimed that they were only looking for a satisfactory outcome from the system and not looking to derive new spatiotemporal information to augment their decisions.
@@
Therefore, to be able to evaluate the efficacy of human-AI decision-making, it is first necessary to ensure that the AI complements the workers’ expertise and augments their decision-making in specific ways. For instance, LEAs shared that it would be meaningful to draw out the boundaries between districts (i.e., insert their hyperlocal contextual knowledge onto the map), look at the relationship between different types of crimes to assess any spatiotemporal relationships, or incorporate information derived from on-the-ground investigations. That is, worker-centered design of AI applications can help us better approach the evaluation of human+AI decisions by ensuring that AI complements workers’ expertise and augments their decisions in measurable ways. This common theme regarding human-AI decision-making has been observed across contexts (e.g., child welfare (Kawakami et al., 2022a; Saxena and Guha, 2023), nursing (Clare Elish and Anne Watkins, 2020), hospitality (Spektor et al., 2023)) where workers drew attention to the nuances of their labor that the AI did not account for and which inadvertently augmented uncertainty and left them performing added labor to repair the disruption in processes caused by the AI system.
@@
Here, ADS such as algorithmic crime mapping could be redesigned to be worker-centered by incorporating LEAs’ feedback. For instance, contextualizing empirical data points with qualitative narratives from police investigations where explanations (i.e., a summary of key points) are generated alongside hotspots to help LEAs assess if new patterns of criminal activity are emerging in neighborhoods and inform collaborative decision-making (See Ehsan et al. (Ehsan et al., 2023) for important practical considerations for designing for explanations where the nature of practice is highly collaborative).
@@ 5.4. Design Guidelines for Algorithmic Crime Mapping that Serves the Public Interest
Our study of a prominent ADS used in the criminal justice system found some comprehensive design principles for creating AI systems for public sector settings. However, as researchers, we must contend with the fact that most of the algorithmic tools used by law enforcement agents in the United States are privately developed with little room for comments from the public. In addition, considering the data being used by such technologies and their disparate use in urban areas for the policing of racial and low-income minorities, we believe that such technologies can lead to significant representational, interpersonal, and social algorithmic harms (Shelby et al., 2023). Therefore, it was essential to explore the community members’ perspectives and gather their feedback. Community members drew upon their lived experiences, the technical participants used their data science knowledge, and the LEAs leveraged their domain expertise as they deliberated while interacting with the application. Below, we provide some design guidelines –
@@
  • Technical participants discussed the elastic nature of data science practice and recommended a more collaborative process for decision-making to ensure accountability and transparency around the practices of crime analysts. They also drew attention to the harmful, data-driven feedback loop created by algorithmic crime mapping.
@@
  • Worker-centered design of AI systems, as specified in the point above, can further make the system more transparent to workers by supporting AI explanations (Ehsan et al., 2021) that are grounded in the nature of practice.
  • It is important to recognize that a target-construct mismatch (Kawakami et al., 2023b) is likely to exist, assess its impact on predicted outcomes, and whether the predicted target still aligns with the practitioners’ decision-making goals.
  • In response to algorithmic unfairness embedded in sociotechnical systems, significant attention has been paid to the biases embedded in datasets (e.g., disproportionate impact on some social groups). By putting the practitioners’ contextual knowledge into action, it is necessary to assess any process-oriented harms (i.e., harm to the decision-making process itself) that may occur (Saxena and Guha, 2023).
@@ 7. Conclusion
In this study, we conducted a mixed-methods investigation in a midwestern, mid-sized city in the United States. The goal of this study was to assess how community stakeholders from different backgrounds interacted with and interpreted algorithmic crime mapping. To achieve this, we designed an in-lab activity where participants were asked to interact with the tool, and subsequently, identify hotspots on the map. Additionally, we sought to gain insights into their perspectives through semi-structured interviews conducted immediately after the activity. We learned that community members found utility in algorithmic crime mapping but were concerned about the incentive structures within police departments themselves that often racially profile and target low-income communities. Technical participants, on the other hand, drew attention to the elastic and subjective nature of data science practice and the need for a more collaborative and interdisciplinary process for decision-making. We learned that the LEAs’ interaction and interpretation of this application was significantly informed by their domain knowledge and they often used it as a means of validating their pre-existing street-level knowledge instead of trying to derive new information from it. We further show that AI systems need to be designed using worker-centered design principles if they are to augment human decisions such that human+AI decisions are an improvement over human-only and AI-only decisions.
@@ Appendix A ESTIMATING THE NUMBER OF HOTSPOTS
Fig 4 shows the distributions of minimum and maximum hotspots grouped by background. For our statistical analysis, we compared user-selected values to actual values for each map type (the solid line represents actual values). We conducted this analysis to ascertain whether people from different backgrounds tend to over or under-estimate hotspots on the crime maps.
@@ Appendix C Assessing mental workload of tasks
The third parameter we considered is performance (PF) which gauges workers’ success and satisfaction. In Figure 8, community members have medium PF scores (10-29), while technical participants and LEAs show somewhat high scores (30-49) (Prabaswari et al., 2019). This suggests greater confidence in results for technical participants and LEAs compared to community members. We explain in Sections 4.1.1 and 4.1.2 that LEAs and technical participants are more familiar with such ADS than community members. Despite technical participants having data science expertise (average PF score of 45.6), LEAs, due to prior experience working with algorithmic crime maps, exhibit similar confidence levels in their outputs.
The fourth parameter we considered is frustration level (FT), which assesses whether the job is frustrating or mentally demeaning such that the participant is left feeling insecure, desperate, offended, or disturbed when doing the task. The average FT is medium (10-29) for community members and low (0-9) for technical participants and LEAs. As we highlight in the qualitative results, community members were more frustrated because they wanted to generate and compare different maps. Notably, LEAs were less frustrated than technical participants, likely because they didn’t generate new maps, relying on the first map and connecting it with their street-level domain knowledge. Through this analysis, we can say that the mental workload regarding the task is not very high which implies it is feasible to use such AI applications as boundary objects that can facilitate deeper conversations about algorithm design and use with a variety of stakeholders.